# LInK: Learning Joint Representations of Design and Performance Spaces through Contrastive Learning for Mechanism Synthesis

### v1

@@ Appendix C Additional Visualizations
Here we will visualize the mechanisms that trace the curves in the 8 test curves and alphabet tests. For each set of tests, we will visualize the kinematics of the entire mechanisms in 2D graph and path figures and provide the optimal assembly of the mechanisms in 3D as well to visualize what the actual assembled mechanism would look like. Note that these solutions are the best perfoming solutions as such they tend to have more joints since higher complexity mechanisms tend to produce more precise solutions.
C.1 Eight Test Curves
@@
Figure 10 shows the mechanisms associated with the solutions found by LInK for each of the 8 test curves. This figure shows the mechanisms and their kinematic in 2D while Figures 11, 12 show the optimal assemblies for each of these mechanisms to avoid collisions while using the fewest layers in the z-axis according to the optimization scheme described in Section 3.4.
@@
Figure 10: This figure shows the best results from LInK on the 8 test curves. As it is visible here LInK has matched the target curves effectively without much challenge. Here we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
@@
C.2 Alphabet Test Curves
The figures in this subsection are similar in nature to the figures in the prior subsection. Again it is important to note the fact that these are the best perfoming solutions which means that even for simpler curves such the letter ‘O’ the produced mechanism is more complex because the more complex mechanisms achieve an exceptionally accurate solution. Note that the LInK approach can be set to be limited any number of joint count as we mentioned in the main text to obtain simpler mechanisms.
Figure 13: This figure shows the best results from LInK on the first 9 alphabet test curves. As it is visible here LInK has matched the target curves effectively without much challenge. Here we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
Figure 14: This figure shows the best results from LInK on the 9-18 alphabet test curves. As it is visible here LInK has matched the target curves effectively without much challenge. Here we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
Figure 15: This figure shows the best results from LInK on the last 8 alphabet test curves. As it is visible here LInK has matched the target curves effectively without much challenge. Here we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
Figure 16: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (A-D).
Figure 17: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (E-H).
Figure 18: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (I-L).
Figure 19: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (M-P).
Figure 20: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (Q-T).
Figure 21: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (U-X).
Figure 22: This figure shows the 3D optimal assemblies for mechanisms displayed in Figure prior figures for the alphabet mechanisms (Y, Z).

### v2

@@ Appendix D Additional Visualizations
Here we will visualize the mechanisms that trace the curves in the 8 test curves and alphabet tests. For each set of tests, we will visualize the kinematics of the entire mechanisms in 2D graph and path figures. We will also provide the optimal assembly of the mechanisms in 3D to visualize what the actual assembled mechanism would look like. Note that these solutions are the best perfoming solutions, as such they tend to have more joints since higher complexity mechanisms tend to produce more precise solutions.
D.1 Eight Test Curves
@@
Figure 13 shows the mechanisms associated with the solutions found by LInK for each of the 8 test curves. This figure shows the mechanisms and their kinematic in 2D while Figures 14, 15 show the optimal assemblies for each of these mechanisms to avoid collisions while using the fewest layers in the z-axis according to the optimization scheme described in Section 3.4.
@@
Figure 13: This figure shows the best results from LInK on the 8 test curves. As it is visible here, LInK has matched the target curves effectively without much challenge. Here, we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
@@
D.2 Alphabet Test Curves
The figures in this subsection are similar in nature to the figures in the prior subsection. Again it is important to acknowledge that these are the best perfoming solutions. Compared to simpler curves –such the letter ‘O’– the produced mechanism is more complex because the more complex mechanisms achieve an exceptionally accurate solution. Note that the LInK approach can be limited to any number of joint counts as mentioned in the main text.
Figure 16: This figure shows the best results from LInK on the first 9 alphabet test curves. As it is visible here, LInK has matched the target curves effectively without much challenge. Here, we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
Figure 17: This figure shows the best results from LInK on the 9-18 alphabet test curves. As it is visible here, LInK has matched the target curves effectively without much challenge. Here, we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
Figure 18: This figure shows the best results from LInK on the last 8 alphabet test curves. As it is visible here, LInK has matched the target curves effectively without much challenge. Here, we visualize the mechanisms in 2D and also plot the paths traced by all the joints in each mechanism.
Figure 19: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (A-D).
Figure 20: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (E-H).
Figure 21: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (I-L).
Figure 22: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (M-P).
Figure 23: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (Q-T).
Figure 24: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (U-X).
Figure 25: This figure shows the 3D optimal assemblies for mechanisms displayed in prior figures for the alphabet mechanisms (Y, Z).